# Implicit neural representations for accurate estimation of the standard model of white matter

### v1

@@ 4.2 Limitations of the approach
The comparison experiments across methods were conducted using Gaussian noise, which differs from the noise characteristics of in vivo magnitude MRI data, typically following Rician or non-central Chi distributions. However, with appropriate preprocessing and ideally the availability of phase data, the noise can be transformed to approximate a Gaussian distribution (Eichner et al., 2015; Koay et al., 2009). This makes the use of Gaussian noise still relevant and consistent with previous work in self-supervised learning for dMRI (Álvaro Planchuelo-Gómez et al., 2024). The Rician noise experiments reveal biases in the parameter estimates by the INR when using MSE, suggesting that this should be taken into consideration. In this work we show the promise of correcting for this bias by using a loss function tailored specifically to Rician noise (Parker et al., 2023).
@@
Additionally, correcting for gradient non-uniformities has a significant impact on the parameter estimates, yet the accuracy remains to be evaluated. Both experiments in section 3.2 and 3.5 show that these corrections result in lower estimates for $D_{e}$ which brings $D_{e}$ more in agreement with previous work showing $D_{i}>D_{e}$ (in gadolinium based contrast experiments (Kunz et al., 2018)) and $D_{i}\sim 2.3~{}\mathrm{\mu m^{2}/ms}$ (in experiments with elaborate acquisition protocols using high diffusion planar tensor encoding (Dhital et al., 2019)). Any further inconsistency with values of $D_{e}$ for the in vivo dataset could be caused by the inability of the acquisition protocol to discriminate solution branches as a high b-shell ($b>5000~{}\mathrm{s/mm^{2}}$) with LTE is lacking (Jelescu et al., 2016; Coelho et al., 2019, 2022). To gain more insight into the degeneracies of the INRs estimation and to enable error quantification, calculating the posterior distribution is required (Consagra et al., 2025; Jallais and Palombo, 2024).
@@ 4.3 Future work
The spatial regularization inherent to INRs in the fitting process can be beneficial for other biophysical models. The presented method focuses on white matter and therefore a white matter segmentation was performed. When sampling outside of the white matter (as shown around the edges in Figure 8), the INR produces inaccurate results. Implementation of gray matter models (Palombo et al., 2020) using INRs avoids the need for fitting solely white matter and can provide whole brain parameter maps.
@@ 4.4 Conclusion
Using INRs to fit the SM provides noise-robust, spatially regularized parameter estimates. FODs of SH orders up to at least eight can be estimated alongside the SM kernel parameters. The unsupervised nature of this approach has advantages over existing (supervised) methods, as it prevents training set bias and allows for explicit correction of gradient non-uniformities, within reasonable estimation times.
@@ 5 Data availability
The MGH Connectome Diffusion Microstructure Dataset (Tian et al., 2022b) is publicly available. The synthetic and in-vivo dataset can be made available upon reasonable request.
@@ 7 Acknowledgements
Appendix A Choice of hyperparameter $\sigma^{2}$
The Fourier Feature (Tancik et al., 2020) encoding provides an additional hyperparameter $\sigma^{2}$ that controls the width of the normal distribution where the encoding frequencies are sampled from. This provides extra flexibility to the user to control the spatial regularization of the representation. Lower $\sigma^{2}$ yields lower frequencies, which capture less detail and noise, while higher $\sigma^{2}$ does the opposite. This implies that downstream tasks which do not require exceptionally fine details could benefit from lower $\sigma^{2}$ to mitigate noise, while higher quality datasets can be modeled with higher $\sigma^{2}$ to maintain contrast in regions with high frequency changes in signal.
To illustrate the effect $\sigma^{2}$ has on the representation, we fit an INR on the synthetic datasets with Gaussian noise SNR 50 and 20 for each $\sigma^{2}\in\{1,2,4,8\}$. The $D_{i}$ parameter map is visualized for every INR in Figure 9.
Figure 9: Parameter maps of $D_{i}$ for the INR fit on the synthetic datasets with SNR 50 and 20 (rows) with varying $\sigma^{2}$ (columns)
At $\sigma^{2}=1$ we see an overly smooth representation, losing some finer detail. These details start to re-appear at $\sigma^{2}=2$. While for SNR 50 the $\sigma^{2}=4$ still provides a good representation, for SNR 20 the INR starts capturing too much of the noise. For $\sigma^{2}=8$ both datasets produce an overly granular parameter map. In general we would recommend starting with $\sigma^{2}=3.5$, and move up or down based on downstream requirements and dataset quality.
Appendix B Noise comparison
Scatter density plots from Fig.10a demonstrate that the most accurate voxel-wise estimates are obtained with NLLS in the noiseless case as expected, where $\rho>0.999$. Only RMSE of $p_{2}$ deviates notably, which is attributed to the lack of constraints on the fODF coefficient values in the NLLS method, leading to outliers. Method 3 also shows high correlation for the kernel ($\rho>0.995$) but shows low accuracy for $p_{2}$. Method 1 shows the lowest accuracy for $f_{i}$ and $D_{p}$. Method 2 performs very accurately for the bulk of the voxels, e.g. less broad scatter density plots than Method 1, but shows high RMSE values for $D_{i}$ and $D_{e}$ due to outliers at low ground truth diffusivity values. Visual inspection of the brain parameter maps in Fig. 10b shows that all fitting methods are capable of producing smooth parameter maps, similar to the ground truth.
Fig.11 shows results for experiment 1 where Gaussian noise is considered with SNR 20. The INR method shows to be noise robust also at higher noise levels and outperforms the other methods. Note the spatial smoothness of the parameter maps estimated by the INR method compared to the other very noisy maps.
Figure 10: Experiment 1 (No Noise): a) Scatter density plots of the ground truth versus the parameter estimations of all methods. Row numbers correspond to method numbering from section 2.8.1. The titles of the subplots indicate $\rho$ and RMSE. b) SM parameter maps corresponding to the results in a. Bottom row shows the ground truth (GT).
Figure 11: Experiment 1 (SNR 50): a) Scatter density plots of the ground truth versus the parameter estimations of all methods. Row numbers correspond to method numbering from section 2.8.1. Title of the subplots indicate $\rho$ and RMSE. b) SM parameter maps corresponding to the results in a. Bottom row shows the ground truth (GT).
Appendix C Rician loss SNR 50
Figure 12: Effect of Rician Loss Likelihood function (SNR = 50). a) Scatter plots for estimation with MSE and Rician loss. b) Brain parameter maps of the predictions from a. Difference maps are with the ground truth parameters.
Appendix D Parameter estimates for higher SH orders
To illustrate the difference in parameter estimates between different $l_{max}$ values, we calculate the median absolute difference between parameter maps. For the synthetic datasets this is calculated between the ground truth parameter maps and the different SH order fits of the INR, the results of which are shown in Table 4. Because there is no ground truth available for the real dataset, we compare the higher SH order estimates with the $l_{max}=2$ fit. These results are shown in Table 5. For both the synthetic and the real dataset there are no notable differences between the different SH order estimates. The noise robustness of the model is shown in the synthetic dataset, where the difference with the ground truth increases only slightly in the SNR = 50 dataset, compared to the noiseless one.
$l_{max}$ Di f De Dp S0
Noiseless 2 0.034 0.006 0.046 0.015 0.728
4 0.034 0.006 0.045 0.015 0.794
6 0.034 0.006 0.044 0.014 0.629
8 0.034 0.006 0.043 0.014 0.547
SNR = 50 2 0.047 0.011 0.051 0.019 1.628
4 0.046 0.010 0.052 0.019 1.696
6 0.046 0.010 0.051 0.019 1.855
8 0.046 0.010 0.051 0.019 1.643
Table 4: Table showing the median absolute differences between the ground truth and two synthetic dataset (noiseless and SNR = 50 gaussian noise) INR fits for different SH orders ($l_{max}$).
$l_{max}$ Di f De Dp S0
Real 2 0.000 0.000 0.000 0.000 0.000
4 0.016 0.005 0.022 0.011 5.725
6 0.019 0.006 0.033 0.015 5.707
8 0.018 0.006 0.033 0.015 4.419
Table 5: Table showing the median absolute differences between the $l_{max}=2$ INR fit for the real dataset and the higher SH order INR fits.

### v2

@@ 4.2 Limitations of the work
The comparison experiments across methods were conducted using Gaussian noise, which differs from the noise characteristics of in vivo magnitude MRI data, typically following Rician or non-central Chi distributions. However, with appropriate preprocessing and ideally the availability of phase data, the noise can be transformed to approximate a Gaussian distribution (Eichner et al., 2015; Koay et al., 2009). This makes the use of Gaussian noise still relevant and consistent with previous work in self-supervised learning for dMRI (Álvaro Planchuelo-Gómez et al., 2024). The Rician noise experiments reveal biases in the parameter estimates by the INR when using MSE, suggesting that this should be taken into consideration. In this work we show the promise of correcting for this bias by using a loss function tailored specifically to Rician noise (Parker et al., 2025).
@@
Additionally, correcting for gradient non-uniformities has a significant impact on the parameter estimates, yet the accuracy remains to be evaluated although comparison to NLLS in combination with gradient non-uniformity correction shows similar results. While changes in shape due to gradient non-uniformities are taken into account ($b_{\Delta}$), a limitation of the current implementation is that it assumes conservation of B-tensor axial symmetry, an assumption that generally does not hold when gradient non-uniformities and non-LTE encodings are considered (Coelho et al., 2025) The exact impact of this approximation would necessitate implementation of SO(3) convolutions and requires further investigation. The experiment in section 3.5 shows that these corrections result in lower estimates for $D_{e}$ which brings $D_{e}$ more in agreement with previous work showing $D_{i}>D_{e}$ (in gadolinium based contrast experiments (Kunz et al., 2018)) and $D_{i}\sim 2.3~\mathrm{\mu m^{2}/ms}$ (in experiments with elaborate acquisition protocols using high diffusion planar tensor encoding (Dhital et al., 2019)). Any further inconsistency with values of $D_{e}$ for the in vivo dataset could be caused by the inability of the acquisition protocol to discriminate solution branches as a high b-shell ($b>5000~\mathrm{s/mm^{2}}$) with LTE is lacking (Jelescu et al., 2016; Coelho et al., 2019, 2022). To gain more insight into the degeneracies of the INRs estimation and to enable error quantification, calculating the posterior distribution is required (Consagra et al., 2025; Jallais and Palombo, 2024).
@@ 4.3 Future work
The spatial regularization inherent to INRs in the fitting process can be beneficial for other biophysical models. The presented method focuses on white matter and therefore a white matter segmentation was performed. When sampling outside of the white matter, the INR produces inaccurate results. Implementation of gray matter models (Palombo et al., 2020) using INRs avoids the need for fitting solely white matter and can provide whole brain parameter maps.
@@ 4.4 Conclusion
Using INRs to fit the SM provides noise-robust, spatially regularized parameter estimates. FODs of SH orders up to at least eight can be estimated alongside the SM kernel parameters. The self-supervised nature of this approach has advantages over existing (supervised) methods, as it prevents training set bias and allows for explicit correction of gradient non-uniformities, within reasonable estimation times.
@@ 5 Data availability
The MGH Connectome Diffusion Microstructure Dataset (Tian et al., 2022b) is publicly available. The synthetic dataset is available at Zenodo ( 10.5281/zenodo.17078893). The in-vivo dataset can be made available upon reasonable request upon signing a data sharing agreement with Cardiff University.
@@ 7 Acknowledgements
This research was funded in whole, or in part, by the Wellcome Trust [215944]. For the purpose of Open Access, the author has applied a CC BY public copyright license to any Author Accepted Manuscript version arising from this submission.
8 Author contributions
Conception: TH, GA, AV, MC, CT; dataset generation: TH, GA, CT; analysis and interpretation of data: TH, GA, EV, AV, MC, CT; drafting manuscript: TH, GA, MC, CT; revising manuscript: TH, GA, EV, AV, MC, CT; funding: AV, MC, CT; supervision: EV, AV, MC, CT. Every author has approved the submitted version, and is personally accountable for their own contributions.
9 Competings interests
The authors declare to have no competing interests.